Patrice alone can paint 4 rooms in 10 hours. If she hires April to help her she can paint 4 rooms in 6 hours. If she lets April paint alone how long will it take her to paint 4 rooms? Thanks you guys are GREAT!
Answer by Earlsdon(6294)

Patrice can do the job in 10 hours, so she can do 1/10 of the job in 1 hour.
Patrice and April together can do the job in 6 hours, so they can do 1/6 of the job in 1 hour.
Therefore: April can do in 1 hour what (Patrice + April) - Patrice can do in 1 hour.
(P + A) - P = 1/6 - 1/10 = 5/30 - 3/30 = 2/30 = 1/15 of the job in 1 hour.
So, April can do 1/15 of the job in 1 hour, so it will take her (15/1) hours to do the entire job.
April will take 15 hours to paint the four rooms.